# Neutrino Oscillation Prospects with a Dual-Baseline Beam from BNL to SNOLAB and SURF

### v1

@@ 6 $\Delta\chi^{2}$ Analysis
In Fig. 9 we show the CP violation sensitivities ($\sigma=\sqrt{\Delta}\chi^{2}$) of the WbLS detectors at SURF and SNO site for a total runtime of 7 yrs. (3.5 yrs. each in $\nu$ and $\bar{\nu}$ modes). The red (green) curves indicate the cases of with (without) energy smearing. The solid curves show the case when we take the full energy interval (0.5180.5-18 GeV) for analysis which considers more than one oscillation maxima. Here in the ideal case of perfect energy reconstruction (solid green), the WbLS detectors at SURF and SNO can reach a sensitivity above $5\sigma$ and $4\sigma$ respectively. Adding a realistic energy resolution for WbLS reduces the sensitivities somewhat, but they still reach above $3.5\sigma$ for SURF, and almost $3\sigma$ for SNO. The dotted curves show the sensitivities with a narrower energy range (2.2182.2-18 GeV for SURF and 0.6180.6-18 GeV for SNO) in order to consider only the first oscillation maximum in the analysis. Using only first oscillation maximum significantly reduces the sensitivities for SURF (Around $3\sigma$ for perfect energy reconstruction and around $2.5\sigma$ with a finite energy smearing) as compared to SNO (around $3.5\sigma$ for no smearing and slightly below $3\sigma$ with a realistic energy smearing). This demonstrates the crucial role played by the second (and higher) oscillation maxima in probing CP violation for longer baselines.
So far, in generating both the solid and the dotted curves, we have marginalised over the less-well known oscillation parameters $\theta_{23}$ and $\Delta m_{31}^{2}$ with some prior uncertainties (see Table 2). Finally, to complete our study we also show the conservative scenario when $\theta_{13}$ is also allowed to vary in an unrestricted manner, - which brings down the CP sensitivities to around $1\sigma$ (dotted curves). Note that, for the dotted curves we have taken the full energy range into account in order to consider multiple oscillation maxima.
In Fig. 10 we show the similar $\Delta\chi^{2}$ results with the assumption that the beam has the full power of 13.2 MW instead of 1 MW. In this optimistic case, the CP sensitivities can be very high, - well above $5\sigma$ even with a finite energy smearing and using only the first oscillation maximum.
@@ 7 Summary and Conclusion
The Electron–Ion Collider (EIC) is an upcoming next-generation particle accelerator designed primarily to study nuclear structure with unprecedented precision by colliding high-energy electron beams with high-energy proton and ion beams. In this work, we have explored how this powerful proton beam can additionally be utilized to generate intense neutrino beams, which can then be studied for oscillation effects using two water-based liquid scintillator (WbLS) detectors—one located at SURF (with a baseline of 2900 km) and another at the SNO site (900 km baseline).
@@
We illustrate the optimized neutrino fluxes at both baselines, along with the corresponding $\text{P}_{\mu e}$ probabilities, neutrino cross sections, and their products, to provide an overview of the expected event spectra. We then discuss the analytical behavior of $\text{P}_{\mu e}$ at different baselines, highlighting the emergence of multiple oscillation maxima. For longer baselines, both the first and second oscillation maxima shift toward higher energies, making them more experimentally accessible. For the 2900 km baseline, the first and second maxima appear around 3.8 GeV and 1.7 GeV, respectively,- energies that lie well within the reach of the EIC-produced neutrino beam. Crucially, probing multiple oscillation maxima significantly enhances the sensitivity to leptonic CP violation. Using analytical expressions and heatmap visualizations, we demonstrate how CP asymmetries vary with $\delta_{\text{CP}}$ across different baselines.
Using GLoBES, we simulate neutrino event spectra for both the $\nu_{e}$ appearance and $\nu_{\mu}$ disappearance channels in a 17 kt (fiducial) WbLS detector placed at SURF (2900 km) and SNO (900 km). We show that the simulated events, both with ideal energy reconstruction and realistic smearing, faithfully reproduce the oscillation features of the underlying probabilities. Finally, we perform a $\Delta\chi^{2}$ analysis to evaluate the CP-violation sensitivity at both baselines. Our results demonstrate the critical role of the second oscillation maximum, particularly for longer baselines in enhancing CP violation discovery potential. For example, with realistic energy smearing, the WbLS detector at SURF (2900 km) achieves a maximum CP sensitivity of about $3.5\sigma$ or higher, while excluding the second oscillation maximum reduces this sensitivity to around $2.5\sigma$. With a near perfect energy reconstruction, the corresponding CP sensitivities are enhanced to around $5\sigma$ and $3\sigma$ respectively.

### v2

@@ 6 $\Delta\chi^{2}$ Analysis
In Fig. 9 we show the CP violation sensitivities ($\sigma=\sqrt{\Delta}\chi^{2}$) of the WbLS detectors at SURF and SNO site for a total runtime of 7 yrs. (3.5 yrs. each in $\nu$ and $\bar{\nu}$ modes). The solid curves show the case when we take the full energy interval (0.5180.5-18 GeV) for analysis which considers more than one oscillation maxima. With a realistic energy resolution the WbLS detectors can achieve sensitivities above $3.5\sigma$ (SURF), and almost $3\sigma$ (SNO). The dotted curves show the sensitivities with a narrower energy range (2.2182.2-18 GeV for SURF and 0.6180.6-18 GeV for SNO) in order to consider only the first oscillation maximum in the analysis. Using only first oscillation maximum significantly reduces the sensitivities for SURF (Around $3\sigma$ for perfect energy reconstruction and around $2.5\sigma$ with a finite energy smearing) as compared to SNO (around $3.5\sigma$ for no smearing and slightly below $3\sigma$ with a realistic energy smearing). This demonstrates the crucial role played by the second (and higher) oscillation maxima in probing CP violation for longer baselines.
So far, in generating both the solid and the dotted curves, we have marginalised over the less-well known oscillation parameters $\theta_{23}$ and $\Delta m_{31}^{2}$ with some prior uncertainties (see Table 2), while keeping the rest of the parameters ($\theta_{12},\theta_{13},\Delta m_{21}^{2}$) fixed to their best fit values (also listed in Table 2). Finally, to complete our study we also show the conservative scenario when $\theta_{13}$ is also marginalized over (in addition to the marginalization of $\theta_{23}$ and $\Delta m_{31}^{2}$), - which brings down the CP sensitivities to around $1\sigma$ (dotted curves). Note that, for the dotted curves we have taken the full energy range into account in order to consider multiple oscillation maxima.
In Fig. 10 we show the similar $\Delta\chi^{2}$ results with the assumption that the beam has the full power of 13.2 MW instead of 1 MW. In this optimistic case, the CP sensitivities can be very high, - well above $5\sigma$ even with using only the first oscillation maximum.
@@ 7 Summary and Conclusion
The Electron-Ion Collider (EIC) is an upcoming next-generation particle accelerator designed primarily to study nuclear structure with unprecedented precision by colliding high-energy electron beams with high-energy proton and ion beams. In this work, we have explored how this powerful proton beam can additionally be utilized to generate intense neutrino beams, which can then be studied for oscillation effects using two water-based liquid scintillator (WbLS) detectors,-one located at SURF (with a baseline of 2900 km) and another at the SNO site (900 km baseline).
@@
We illustrate the optimized neutrino fluxes at both baselines, along with the corresponding $\text{P}_{\mu e}$ probabilities, neutrino cross sections, and their products, to provide an overview of the expected event spectra. We then discuss the analytical behaviour of $\text{P}_{\mu e}$ at different baselines, highlighting the emergence of multiple oscillation maxima. For longer baselines, both the first and second oscillation maxima shift toward higher energies, making them more experimentally accessible. For the 2900 km baseline, the first and second maxima appear around 3.8 GeV and 1.7 GeV, respectively,- energies that lie well within the reach of the EIC-produced neutrino beam. Crucially, probing multiple oscillation maxima significantly enhances the sensitivity to leptonic CP violation. Using analytical expressions and heatmap visualizations, we demonstrate how CP asymmetries vary with $\delta_{\text{CP}}$ across different baselines.
Using GLoBES, we simulate neutrino event spectra for both the $\nu_{e}$ appearance and $\nu_{\mu}$ disappearance channels in a 17 kt (fiducial) WbLS detector placed at SURF (2900 km) and SNO (900 km). We show that the simulated events, both with ideal energy reconstruction and realistic smearing, faithfully reproduce the oscillation features of the underlying probabilities. Finally, we perform a $\Delta\chi^{2}$ analysis to evaluate the CP-violation sensitivities at both baselines. Our results demonstrate the critical role of the second oscillation maximum, particularly for longer baselines in enhancing CP violation discovery potential. For example, with realistic energy smearing, the WbLS detector at SURF (2900 km) achieves a maximum CP sensitivity of about $3.5\sigma$ or higher, while excluding the second oscillation maximum reduces this sensitivity to around $2.5\sigma$.
It should be noted that this study uses simplified systematics and neglects backgrounds; a full simulation including these effects is reserved for future work.